# Distributional Scaling Laws for Emergent Capabilities

### v2

@@ 3.3 Bimodality Emerges Abruptly
In the previous section, we found that bimodal random distributions cause performance breakthroughs when scaling up a single training run. We also wish to analyze how the distribution itself evolves across scales; as shown in Figure 2, the performance distribution starts as unimodal at the smallest scale, where all models are unable to length generalize. Larger scales yield a bimodal distribution where most of the probability mass is ultimately placed on successful length generalizing runs. A priori, there could be a smooth evolution between these two distributions in which probability mass from failing runs gradually shifts towards higher performance metrics, eventually splitting into a clear separate cluster. We now show, to the contrary, that the shift from unimodal to bimodal is abrupt and instantly polarized into low and high clusters.
To track the evolution of the performance distribution across scales, we plot the Wasserstein-L2 distance of each distribution relative to that of the final, largest model scale when separately fixing width and depth. In Figure 4 we see that there is a sharp decrease in the W2 distance for scaling with a fixed width, quantifying the sudden appearance of highly successful runs when model depth reaches 3 layers or model width reaches 2 heads. These sudden changes identify the moment when a new capability is unlocked, as the distribution transitions abruptly from unimodal on one extreme to bimodal at both extremes. We posit that this transition marks the minimum capacity required to learn the task. We also mark the point in each trend where the mode in Figure 3 (top left) increases sharply. We posit that it may be misleading to draw conclusions about minimal model capacity on a specific task using single runs at each scale, whereas distributional metrics correctly predict the sudden appearance of probability mass placed on successful runs.
@@ 3.4 Count: An Instance of U-shaped Scaling
We next consider the counting task (Section 2.1), which also exhibits bimodally-distributed performance (see Figure 5) but ultimately yields a very different scaling effect: a U-shaped curve. The summary statistics in Figure 6 reveal this peculiar phenomenon, particularly in the fixed-depth trendline. This U-shaped trend in mean accuracy is similar to that observed on certain natural language tasks (Wei et al., 2023). Furthermore, this phenomenon is not simply a result of the choice of summary statistic, as the same curve describes the evolution of the distributions as a whole when measuring their W2 distance in Figure 7.
U-shaped scaling has been observed in LLMs, but its causes are not currently well-understood. When the Inverse Scaling Prize (McKenzie et al., 2022) solicited tasks which exhibit inverse scaling trends—performance decreases at scale—for large models, Wei et al. (2023) revealed that the majority of awarded tasks actually exhibit U-shaped scaling after considering even larger models. Treating the counting task as a concrete instance of U-shaped scaling at small model scales, we find that this unusual trend is still underscored by monotonic continuous changes in the performance distribution. Indeed, Figure 6 (bottom right) shows that although the trend in the mean across all runs is U-shaped curve, the mean of the “successful” runs—those achieving at least 50% accuracy—still improves monotonically when increasing width. The observation of inverse scaling is, instead, due to changes in the probability of success (bottom left). Even when inverse scaling is in effect across a performance distribution, the performance of successful runs may exhibit more conventional responses to scale.
Figure 6: As in Figure 3, we capture various summary statistics for test length 60 on the count task. For fixed width (scaling depth), the trends across overall mode, overall mean, fraction, and mean of successful (above 50% accuracy) runs are similar to the reverse order addition case. The trend for fixed depth (scaling width) exhibits a U-shaped curve emerge for the mean across all runs; however, the mean of successful runs still exhibits continuous improvement.
Figure 7: Wasserstein-L2 distance relative to final model scale when scaling depth and width for the count task. When fixing depth and scaling width, we observe that the W2 metric has an upside-down U-shape at an intermediate range of model scales.
@@ 3.5 Is Bimodality a Mirage?
Metrics with hard thresholds and discontinuities can artificially induce breakthrough behavior (Schaeffer et al., 2024); conversely, continuous metrics can make apparently emergent curves into smooth curves (Srivastava et al., 2023). We must be particularly cautious about claiming emergence when requiring model outputs to exactly match a target string, as we do for both synthetic tasks. Are our case studies artifacts of thresholding effects? In other words, do their bimodal distributions become unimodal under continuous metrics? To assess the role of the exact match metric in creating apparent bimodality, we instead consider a continuous equivalent to the exact match metric: the minimum probability assigned to any individual token. If the model fails to assign its highest probability to the correct token for any position in the sequence, it receives a score of zero under the exact match metric, even if the correct token was still highly ranked. Likewise, if the model successfully assigns maximum probability to the correct token, exact match considers it to be accurate whether that probability was high or low. Our continuous metric reports the worst mistake made by the model, i.e., the minimum probability of any correct token in the sequence.
@@
In Figure 8, we plot a histogram for this continuous metric averaged across all samples. The distribution of this metric across random seeds is often still clustered; its bimodality is not due to thresholding alone. We therefore confirm that emergent capabilities exhibit bimodal performance distributions even when using a continuous performance metric.
(a) Fixing network depth to 4 layers.
(b) Fixing network width to 256 hidden dimension.
Figure 8: Histograms of the probabilities assigned to the correct token, calculated by taking the minimum probability within each sequence and averaging across all samples. These histograms are generated for a test length of $60$ on the count task, scaling the parameter count by increasing width with a fixed model depth (top) and by increasing depth with a fixed model width (bottom). Random variation still leads to bimodal performance distributions, even using this continuous performance metric.
@@
Figure 9: Violin plots of last-layer reinitialized Qwen2.5 models on the MMLU benchmark. Left Models are trained on different mixtures of the C4 news subset and the MMLU auxiliary training set. With insufficient MMLU training data, Qwen2.5-0.5B fails to follow the multiple-choice format, resulting in trivial performance. As the MMLU data ratio increases, models trained on different seeds exhibit a bimodal distribution in performance. With enough MMLU examples, all models score above the random baseline. Right: For a fixed data mix, the smaller model (Qwen2.5-0.5B) is constrained by its capacity and shows a bimodal distribution in MMLU performance. In contrast, the larger model (Qwen2.5-1.5B) consistently learns to answer multiple-choice questions requiring natural language understanding. However, our training data for continued pretraining lacks diversity and therefore, models trained on them do not fully recover the multiple choice capability of the original model.
@@ 4 Natural Language Results
Having documented the bimodal variation of emergent capabilities in small synthetic settings, we turn to large language models (LLMs). We focus on the MMLU dataset, where high performance emerges after the LLM learns the multiple choice format (Srivastava et al., 2023; Hu & Frank, 2024). To avoid the expense of repeatedly training large-scale models from scratch, we simulate independent runs by reinitializing the upper layer of pretrained LLMs before continuing to train them.
@@ 4.1 Emergence Across Data Compositions
In Figure 9 (left), when trained on a data mix containing only 5% MMLU examples and 95% random sequences from C4 (Raffel et al., 2023), most models (out of 20 seeds) achieve near 0% performance on the MMLU test set. According to the existing literature, this failure stems from a failure to process the multiple-choice format (Hu & Frank, 2024). As the proportion of MMLU samples increases, bimodality emerges when MMLU composes 10% of the training data, where models form two distinct performance clusters. In one cluster, continually pretrained models still fail to follow the correct format, while in the other, models achieve performance at or above the random guess baseline of 25%. This second cluster contains models that consistently respond with valid multiple choice options; those that outperform the random baseline have even learned to compose the multiple choice format with their world knowledge when selecting an answer. Finally, with sufficient data ($>20\%$ MMLU), models across all 20 seeds consistently outperform the 25% random baseline. [^2] Note that Qwen2.5-0.5B base model can already following instructions 0-shot without further finetuning, achieving 39.5% performance on MMLU. By re-intializing the last layer, we removed the base model’s MMLU ability. Our continued training dataset, a mix of C4 news data and the MMLU auxiliary training data, is insufficient to recover the full capability of the the base model. Appendix D speculates further on the effect of continued training with this pre-existing capability.
@@ 4.2 Emergence Across Model Scales
Next, we show that, as in the synthetic setting, LLM model size affects the formation of performance clusters. Fixing the dataset to a mixture with 10% MMLU auxiliary training samples, we continually pretrain Qwen2.5-1.5B on the same 20 seeds and report MMLU test accuracy in Figure 9 (right). While models built on the smaller Qwen2.5-0.5B form two distinct performance clusters, those built on the larger Qwen2.5-1.5B consistently outperform the random baseline. We conclude that larger scale models can reliably acquire MMLU capability when trained on the same dataset that produces highly bimodal variation at smaller scales. This conclusion recalls the literature on scaling laws, which argues that smaller models require more training examples to match the performance of larger models (Rosenfeld et al., 2019; Kaplan et al., 2020).
@@ 5 Related Works
Random variation: Even after controlling for critical hyperparameters such as learning rate, model performance remains sensitive to stochastic aspects of the training process—specifically, random initialization and the order of training examples. Prior studies have documented consistent performance differences across various stress test sets (D’Amour et al., 2022; Naik et al., 2018) and observed that these differences persist throughout training, not just at the final checkpoint (Zhou et al., 2020). Dodge et al. (2020) compared the impacts of weight initialization and data ordering, concluding that both contribute equally to variations in out-of-sample performance. Additionally, Sellam et al. (2022) demonstrated that different BERT models independently pre-training from scratch can differ in downstream task performance and social biases. Existing work has also found model runs can cluster in out-of-distribution behavior (Juneja et al., 2023) and in training dynamics (Qin et al., 2024; Hu et al., 2023), hinting at multimodal variation. Our work connects known these random clustering effects to the phenomenon of emergence at scale.
@@
Length generalization and compositionality with transformers: Several studies have highlighted the challenge of length generalization in transformers (Anil et al., 2022; Deletang et al., 2023; Gontier et al., 2020; Hupkes et al., 2020; Schwarzschild et al., 2021; Zhang et al., 2022). Various remedies have been proposed to address length generalization failures. Some methods focus on alternative positional encodings (Shaw et al., 2018; Press et al., 2022; Su et al., 2024; Kazemnejad et al., 2024; Jelassi et al., 2024). Others modify the dataset format, either adding scratchpad or Chain-of-Thought formats (Anil et al., 2022) or incorporating padding and index hints for specific arithmetic tasks (Jelassi et al., 2023; Zhou et al., 2024a). Regarding random variation, Zhou et al. (2024a) and Zhou et al. (2024b) provide evidence of variability in length generalization across random seeds, which we further investigate and analyze across a range of model scales.
@@
Emergent abilities of LLMs: In large language models, emergent abilities are behaviors that arise unexpectedly as models are scaled up in size or trained on larger datasets (Hestness et al., 2017; Rosenfeld et al., 2019; Brown et al., 2020; Kaplan et al., 2020). These abilities are characterized by unpredictable and abrupt performance improvements on specific benchmarks at certain scales (Wei et al., 2022; Ganguli et al., 2022; Srivastava et al., 2023). Understanding the conditions and mechanisms underlying emergence is a key area of research. Recent studies suggest that emergent abilities may stem from the choice of evaluation metrics rather than fundamental changes in model behavior with increased scale (Schaeffer et al., 2024). Nonetheless, some breakthrough capabilities remain emergent. One direction to better understand the mechanisms underlying emergence is through studying language models trained on algorithmic tasks exhibiting similar behavior; for instance, Gopalani et al. (2024) show that BERT models learn the low-rank matrix completion problem with a sudden drop in the loss before interpreting components of the model before and after this transition. Snell et al. (2024) found that some scales exhibit earlier emergence if finetuned explicitly on an emergent task, suggesting that smaller models may have the capacity for that task but are limited by its scarcity in the training corpus. In a similar vein, we show that emergent capabilities can arise from multimodal random variation using synthetic length generalization tasks as a case study.
@@
Depth versus Width Scaling: Despite scaling laws offering a smooth extrapolation for model performance, downstream performance can vary depending on architecture shape and not just model size (Tay et al., 2022). For compositional tasks, deeper models often generalize better up to a certain point, but for a fixed compute budget, it may be more advantageous to train a shallower, wider model (Petty et al., 2024). Various works have proposed explanations for the role of width versus depth in scaling behavior; for instance, Edelman et al. (2024) show that increasing network width offers more ‘parallel queries’ over randomized subnetworks which learn sparse features more efficiently. Levine et al. (2020) use a border rank argument to establish a width-dependent depth threshold, beyond which additional depth yields diminishing returns. In this work, we specifically investigate how independently scaling width and depth influences the random variation distribution in compositional tasks.
@@ 6 Discussion and Conclusions
Our work explores the evolution of random variation in model performance across scales, bringing a more nuanced perspective on emergent capabilities than the conventional approach of plotting a single model run per scale. Previous work has already documented the variability of length generalization across random seeds (Zhou et al., 2024b, a); in general, out-of-distribution behavior like compositional rules (McCoy et al., 2019) or associative biases (Sellam et al., 2022) often exhibit extreme variation compared to in-distribution performance. We attribute a capability’s emergence to its underlying bimodal random distribution, which is also documented in text classifier performance metrics (Juneja et al., 2023) and even in the timing of generalization breakthroughs during training (Hu et al., 2023). The mode of these performance distributions displays a sharp improvement at a certain model scale, similar to emergent capabilities in the literature, but we attribute these sudden jumps to gradual improvements in the random distribution. Furthermore, we show that bimodality can emerge before the mode—or most seeds—exhibit a breakthrough (Section 3.3); the transition from the initial unimodal distribution to a bimodal one is sudden.
In the wider scaling laws literature, which often focuses on aggregated loss, parameters added by increasing depth and width are often treated interchangeably in their contribution to total compute. By separately considering depth and width, our results suggest that emergent phenomena may be sensitive to tradeoffs between architectural hyperparameters. In particular, we document a regime in which increasing width damages model performance but increasing depth improves model performance. In a research environment increasingly concerned with compositional and breakthrough capabilities, our findings should inspire further study of how these tasks respond to architectural hyperparameter tradeoffs. Our synthetic case study on U-shaped scaling can also enable future research into the origins of inverse scaling trends.
Future work should rigorously verify these findings at larger scales and across more diverse tasks. We especially hope future research can confirm that emergent benchmarks are bimodal when pretraining from scratch across different seeds.
Impact Statement
This work challenges conventional views on emergent capabilities in large language models by demonstrating that performance breakthroughs can be explained by multimodal distributions across training seeds rather than sharp phase transitions. By highlighting the role of random variation in scaling laws, our findings provide a new perspective on model evaluation and reproducibility, emphasizing the need for robust statistical analysis in benchmarking. This research has broad implications for interpreting model capabilities and the responsible development of AI systems.

### v3

@@ 3.3 Bimodality Emerges Abruptly
The previous section showed that bimodal performance distributions produce emergent scaling curves for individual seeds. We next analyze how the distribution itself evolves across scales. As shown in Figure 2, the performance distribution starts as unimodal at the smallest scale, where no models can length generalize. Larger scales yield a bimodal distribution where most of the probability mass is ultimately placed on successful length generalizing runs. A priori, there could be a smooth evolution between these two distributions in which probability mass from failing runs gradually shifts towards higher performance metrics, eventually splitting into a clear separate cluster. We find instead that the shift from unimodal to bimodal is abrupt and instantly polarized into low and high clusters. In Appendix Figure 10 we show that models fail to generalize when fixing depth to be one layer even when scaling width (despite having perfect in-distribution accuracy).
To track the evolution of the performance distribution across scales, we plot the Wasserstein-L2 distance of each distribution relative to that of the final, largest model scale when separately fixing width and depth. In Figure 4(a) we see that there is a sharp decrease in the W2 distance for scaling with a fixed width, quantifying the sudden appearance of highly successful runs when model depth reaches 3 layers or model width reaches 2 heads. These sudden changes identify the moment when a new capability is unlocked, as the distribution transitions abruptly from unimodal on one extreme to bimodal at both extremes. We posit that this transition marks the minimum capacity required to learn the task. We also mark the point in each trend where the mode in Figure 3 (top left) increases sharply. We posit that it may be misleading to draw conclusions about minimal model capacity on a specific task using single runs at each scale, whereas distributional metrics correctly predict the sudden appearance of probability mass placed on successful runs.
@@ 3.4 Count: An Instance of U-shaped Scaling
We next consider the counting task (Section 2.1), which also exhibits bimodally-distributed performance (see Appendix Figure 9) but ultimately yields a very different scaling effect: a U-shaped curve. Figure 5 reveals this peculiar phenomenon in the mean accuracy scaling curve, when holding depth fixed. This curve is not simply a result of the summary statistic chosen, as it is mirrored by the evolution of the distributions as a whole according to their W2 distance in Figure 4(b).
U-shaped scaling has been observed in LMs, but its causes are not currently well-understood. When the Inverse Scaling Prize [28] solicited tasks which exhibit inverse scaling trends—performance decreases at scale—for large models, Wei et al. [46] revealed that the majority of awarded tasks actually exhibit U-shaped scaling after considering even larger models. Treating the counting task as a concrete instance of U-shaped scaling at small model scales, we find that this unusual trend is still underscored by monotonic continuous changes in the performance distribution. Indeed, Figure 5 (bottom right) shows that although the trend in the mean across all runs is U-shaped curve, the mean of the “successful” runs—those achieving at least 50% accuracy—still improves monotonically when increasing width. The observation of inverse scaling is, instead, due to changes in the probability of success (bottom left). Even when inverse scaling is in effect across a performance distribution, the performance of successful runs may exhibit more conventional responses to scale.
Figure 5: Summary statistics for length generalization (count task). Models tested at length 60, scaling width (with fixed depth 4) and depth (with fixed width 4) independently. The trend for fixed depth (scaling width) exhibits a U-shaped curve for the mean across all runs 5; however, the mean of successful runs (defined as >50% EM accuracy) 5 still exhibits continuous improvement.
@@ 3.5 Is Bimodality a Mirage?
Metrics with hard thresholds and discontinuities can artificially induce breakthrough behavior [36]; conversely, continuous metrics can make apparently emergent curves into smooth curves [41]. We must be particularly cautious about claiming emergence when requiring model outputs to exactly match a target string, as we do for both synthetic tasks. Are our case studies artifacts of thresholding effects? In other words, do their bimodal distributions become unimodal under continuous metrics? To assess the role of the exact match metric in creating apparent bimodality, we instead consider a continuous equivalent to the exact match metric: the minimum probability assigned to any individual token. If the model fails to assign its highest probability to the correct token for any position in the sequence, it receives a score of zero under the exact match metric, even if the correct token was still highly ranked. Likewise, if the model successfully assigns maximum probability to the correct token, exact match considers it to be accurate whether that probability was high or low. We now consider a continuous version of EM accuracy that reports the worst mistake made by the model. This continuous error score for model $f$ on dataset $X$ is based on the per-token loss $L(f(x_{0\ldots i-1}),x_{i})$:
$\textrm{error}(f,X)=\frac{1}{|X|}\sum_{x\in X}\max_{i<|x|}L(f(x_{0\ldots i-1}),x_{i})$ (1)
@@
In Figure 6, we plot this continuous metric for addition models on the length generalization test dataset. The distribution of this metric across random seeds is often still clustered. We therefore confirm that emergent capabilities exhibit bimodal performance distributions even when using a continuous performance metric; their bimodality is not due to thresholding alone. Further discussion around continuous metric plots for count and histograms for probability-based metrics are in Appendix D.
(a) Fixing network depth to 6 layers while scaling up width.
(b) Fixing network width to 8 heads per layer while scaling up depth.
Figure 6: Random variation in continuous length generalization error (addition task). Kernel Density Estimation (KDE) of our loss-based error metric (Equation 1) across model runs. At scales where the EM accuracy distribution is bimodal, the distribution remains bimodal even when using a continuous metric.
(a) Qwen2.5-0.5B trained on different mixes.
(b) Qwen2.5 models trained on 10% MMLU data.
@@
Figure 7: Performance on MMLU. All models have their last layer reinitialized randomly and then are trained on a mix of C4 news and MMLU, with varying proportions of MMLU. Left: With insufficient MMLU training data, each model fails to follow the multiple-choice format, resulting in trivial performance. As more MMLU data is included, different seeds lead to bimodally distributed performance. With enough MMLU data, all models score above the random baseline. Right: For a fixed data mix, the smaller model has bimodally distributed performance. In contrast, the larger model consistently relearns to answer multiple-choice questions, though not at its original base performance.
@@ 4 Natural Language Results
After connecting emergence with bimodality in small synthetic settings, we turn to LMs. We focus on the MMLU dataset, where high performance emerges after the LM learns the multiple choice format [41, 13]. To avoid the expense of repeatedly training large-scale models from scratch, we simulate independent runs by reinitializing the upper layer of pretrained LMs before continuing to train them. Because Qwen base models perform well for their scale (39.5% for Qwen2.5-0.5B), these experiments effectively require models to re-discover MMLU capabilities after their removal. Our continued training dataset, a mix of C4 news data and the MMLU auxiliary training data, is insufficient to recover the full MMLU capability of the the base model.
@@ 4.1 Emergence Across Data Compositions
In Figure 7 (left), when trained on a data mix containing only 5% MMLU examples and 95% random sequences from C4 [34], most models (out of 20 seeds) achieve near 0% performance on the MMLU test set. According to the existing literature, this failure stems from a failure to process the multiple-choice format [13]. As the proportion of MMLU samples increases, bimodality emerges when MMLU composes 10% of the training data, where models form two distinct performance clusters. In one cluster, continually pretrained models still fail to follow the correct format, while in the other, models achieve performance at or above the random guess baseline of 25%. This second cluster contains models that consistently respond with valid multiple choice options; those that outperform the random baseline have even learned to compose the multiple choice format with their world knowledge when selecting an answer. Finally, with sufficient data ($>20\%$ MMLU), models across all 20 seeds consistently outperform the 25% random baseline.
@@ 4.2 Emergence Across Model Scales
Next, we show that, as in the synthetic setting, LM model size affects the formation of performance clusters. Fixing the dataset to a mixture with 10% MMLU auxiliary training samples, we continually pretrain Qwen2.5-1.5B on the same 20 seeds and report MMLU test accuracy in Figure 7 (right). While models built on the smaller Qwen2.5-0.5B form two distinct performance clusters, those built on the larger Qwen2.5-1.5B consistently outperform the random baseline. We conclude that larger scale models can reliably acquire MMLU capability when trained on the same dataset that produces highly bimodal variation at smaller scales. This conclusion recalls the literature on scaling laws, which argues that smaller models require more training examples to match the performance of larger models [35, 21]. In Appendix E.1, we show similar results in the Commonsense QA [43] task.
@@ 5 Related Works
Random variation: Model performance can be sensitive to stochastic aspects of the training process like random initialization and training data order. Prior studies have documented performance differences across various stress test sets [4, 30]. Such differences persist throughout training, not just at the final checkpoint [50]. Dodge et al. [6] compared the impacts of weight initialization and data ordering, concluding that both contribute equally to variation in performance. Additionally, Sellam et al. [38] demonstrated that LMs independently pre-trained from scratch can differ in downstream task performance and social biases. Existing work has also found model runs can cluster in out-of-distribution behavior [19] and in training dynamics [33, 14], hinting at multimodal variation. Our work connects these random clustering effects to the phenomenon of emergence at scale.
@@
Length generalization and compositionality with transformers: Transformer models cannot always generalize to longer sequences than they are trained on [1, 5, 9, 15, 37, 48]. Various proposals have, consequently, aimed to improve length generalization. Some methods focus on alternative positional encodings [39, 32, 42, 22, 17]. Others modify the dataset format by adding scratchpad or Chain-of-Thought [1] or by incorporating padding and index hints [16, 49]. Regarding random variation, Zhou et al. [49] and Zhou et al. [51] provide evidence of variability in length generalization across random seeds, which we further investigate and analyze across a range of model scales.
@@
Emergent abilities of LMs: In large language models, emergent abilities are behaviors that arise unexpectedly as models are scaled up in size or trained on larger datasets [12, 35, 2, 21]. These abilities are characterized by unpredictable and abrupt performance improvements on specific benchmarks at certain scales [45, 8, 41]. Understanding the conditions and mechanisms underlying emergence is a key area of research. Although recent studies suggest that some breakthroughs may stem from the choice of evaluation metrics rather than fundamental changes in model behavior with increased scale [41, 36], some breakthrough capabilities remain emergent. Snell et al. [40] found that some scales exhibit earlier emergence if finetuned explicitly on an emergent task, suggesting that smaller models may have the capacity for that task but are limited by its scarcity in the training corpus. Similarly, we show that emergent capabilities can arise from multimodal random variation using synthetic length generalization tasks as a case study.
@@
Depth versus Width Scaling: Although scaling laws offer a smooth extrapolation for model performance, downstream performance can vary depending on architecture shape and not just model size [44]. For compositional tasks, deeper models often generalize better up to a certain point, but for a fixed compute budget, it may be more advantageous to train a shallower, wider model [31]. Various works have proposed explanations for the role of width versus depth in scaling behavior; for instance, Edelman et al. [7] show that increasing network width offers more ‘parallel queries’ over randomized subnetworks which learn sparse features more efficiently. Levine et al. [24] use a border rank argument to establish a width-dependent depth threshold, beyond which additional depth yields diminishing returns. In this work, we specifically investigate how independently scaling width and depth influences the random variation distribution in compositional tasks.
@@ 6 Discussion and Conclusions
Our work explores the evolution of random variation in model performance across scales, bringing a more nuanced perspective on emergent capabilities than the conventional approach of plotting a single model run per scale. Previous work has already documented the variability of length generalization across random seeds [51, 49]; in general, out-of-distribution behavior like compositional rules [27] or associative biases [38] often exhibit extreme variation compared to in-distribution performance. We attribute a capability’s emergence to its underlying bimodal random distribution, which is also documented in text classifier performance metrics [19] and even in the timing of generalization breakthroughs during training [14]. The mode of these performance distributions displays a sharp improvement at a certain model scale, similar to emergent capabilities in the literature, but we attribute these sudden jumps to gradual improvements in the random distribution. Furthermore, we show that bimodality can emerge before the mode—or most seeds—exhibit a breakthrough (Section 3.3); the transition from the initial unimodal distribution to a bimodal one is sudden.
In the wider scaling laws literature, which often focuses on aggregated loss, parameters added by increasing depth and width are often treated interchangeably in their contribution to total compute. By separately considering depth and width, our results suggest that emergent phenomena may be sensitive to tradeoffs between architectural hyperparameters. In particular, we document a regime in which increasing width damages model performance but increasing depth improves model performance. In a research community increasingly concerned with emergent capabilities, our findings should inspire further study of how these tasks respond to architectural hyperparameter tradeoffs.